# Reinforcement learning for anisotropic p-adaptation and error estimation in high-order solvers

### v1

@@ 4.2 Comparison of p-adaptation approaches and accuracy in lift and drag for a cylinder at Reynolds 100
Once we have tested that the performance of the p-adaptation strategy is appropriate and that we can trust, in general, the error estimation, we increase the complexity of the problem to validate different aspects of the RL p-adaptation approach. In particular, to improve the quality of the designed test cases and quantitatively measure the performance of the p-adaptation strategy, we compare our novel approach with a p-adaptation sensor based on the original sensor developed by Persson and Peraire [19]. This sensor compares the solution inside each element using a different number of high-order modes, providing a spectral decay indicator that can be used as an estimation of the smoothness of the solution. Although it was first developed for shock-capturing purposes, based on this sensor some p-adaptation approaches have been successfully used in DG solvers [17]. However, the available strategies based on this sensor are very limited as they have only been used to increase the polynomial order (not to decrease it) and they are cannot handle anisotropic adaptation. Therefore, we have slightly improved the approaches based on this sensor to add these two functionalities; in this way, we can perform a fair comparison between our novel approach and the state-of-the-art sensor.
The problem to be solved is the laminar, but unsteady, flow field at $Re=100$ around a cylinder at $M=0.15$. The mesh used for this problem is shown in Figure 10. This coarse mesh has only 1282 elements on purpose, again to highlight the ability of the RL agent to increase the polynomial where required. However, the elements are slightly smaller downstream to capture the vortex shedding.
@@
Figure 10: Representation of the mesh used for the simulation of the flow around a cylinder at $Re=100$.
For this problem, a dynamic p-adaptation is required, as the solution is unsteady. A snapshot of the average polynomial order in each element using both approaches (RL and Persson and Peraire-based sensor), once the flow field is stabilized, is represented in Figure 11. At first glance, both approaches are similar, as they increase the polynomial near the cylinder and in the wake, where the vortex shedding is present. However, the adaptation performed by our RL agent is more efficient as it increases the polynomial order only where required, while the Persson and Peraire-based approach increases the polynomial order in some regions upstream and also downstream far from the wake, where the flow field is almost steady.
It is very important to mention that, on the one hand, the Persson and Peraire-based approach relies on a threshold to decide when to increase and decrease the polynomial order and this threshold is problem-dependent. It took several attempts of trial and error, using different values of the aforementioned threshold, to achieve the polynomial adaptation shown in Figure 11(b). On the other hand, the parameters of our RL agent are decided beforehand, before the training, and their dependency with each specific problem is small. Therefore, our agent does not require problem-specific tuning, which is an important advantage of the proposed methodology, when compared to other state-of-the-art p-adaptation strategies.
For completeness, a snapshot of the velocity field is represented in Figure 12. The vortex shedding is well captured, as expected. Finally, we report the accuracy of the p-adaptation approach in Table 3. A comparison between the average drag, the maximum lift and the Strouhal number is considered. The results obtained with the proposed methodology are very close to those of other sources.
@@
Figure 11: Comparison of the p-adaptation between a) the RL agent strategy and b) the improved Persson and Peraire sensor for a simulation of the flow around a cylinder at $Re=100$. A snapshot of the average polynomial order in each element of the mesh is illustrated.
@@
Figure 12: Snapshot of the velocity field for the simulation of the flow around a cylinder at $Re=100$ using the RL p-adaptation agent. Both, a) the horizontal velocity and b) the vertical velocity are shown.
@@ 4.3 Laminar - turbulent transition for the Taylor Green Vortex problem at Reynolds 1600
The flow field is continuously evolving and it has been widely reported that under-resolution without special numerical treatment creates poor results [58, 59, 60, 61], therefore a dynamic p-adaptation is an excellent option to accurately solve the problem while maintaining the cost at minimum. A p-adaptation is performed at every $0.05$ non-dimensional time units. The non-dimensional velocity magnitude and the polynomial adaptation for different timestamps are shown in Figure 13. Each snapshot shows how the p-adaptation algorithm is able to capture velocity fluctuations by increasing the polynomial order locally. As time progresses, the flow transitions to turbulent, and hence the average polynomial order in the whole domain increases. This phenomenon is highlighted in Figure 14(a), where it is clear how the polynomial order is progressively increased, almost reaching the maximum polynomial order $p_{\max}=6$ after 20 non-dimensional time units (fully turbulent regime).
Finally, in Figure 14(b) we show the evolution of the kinetic energy rate over time for different solutions: two uniform polynomial orders $p=3$ and $p=6$, the RL p-adaptation and a DNS solution of [62]. On the one hand, the coarse $p=3$ solution presents large deviations from the reference DNS solution at the transitional times $t/t_{c}\in[8,14]$. On the other hand, both the p-adapted and the uniform $p=6$ solution are very close to the reference DNS simulation, with both matching almost perfectly. Despite these similar results between the $p=6$ and the p-adapted solutions, the simulation with dynamic p-adaptation was much faster, as described in Table 4. It is shown that the novel p-adaptation strategy is accurate in the presence of turbulence while significantly reducing the computational cost, with a cost reduction close to the 60% of savings.
@@ 4.4 Simulation of the wind turbine DTU 10MW
Finally, once the reinforcement learning agent has been tested in a wide variety of scenarios and problems, we apply this methodology to solve the flow field around a wind turbine. The objective behind this problem is to demonstrate that the proposed p-adaptation methodology can be used in complex scenarios without any previous knowledge (during training) regarding the simulation.
We focus on the DTU 10 MW wind turbine, whose blades of the wind turbine are modeled using an actuator line [63, 64], while the nacelle and pylon have been simulated using an immersed boundary approach [65, 49]. Furthermore, a no-slip boundary condition has been set at the bottom of the domain, while free-slip boundary conditions have been applied at the top and at the side faces of the domain. Each of these elements adds an additional level of complexity to the p-adaptation algorithm:
@@
    Behind the rotor, in the wake of the wind turbine, the flow is fully turbulent and high polynomials are required. An accurate computation of this region is essential to obtain accurate predictions of the power generated by the wind turbine.
@@
    An LES model is used to solve under-resolved scales. In this case, we use the Vreman model [66].
All these elements must be taken into account to obtain an accurate solution. In our case, we use a coarse and almost uniform Cartesian h-mesh with $294492$ hexahedral elements, and the RL agent will select the appropriate polynomial inside each element. We limit the maximum polynomial order to $p=4$ due to computational constraints. Again, it is exactly the same RL agent that was used in all the previous cases and it does not have any knowledge in relation to this specific problem. Despite this, we expect to obtain an appropriate polynomial adaptation based on the values at the Gauss nodes (the state) that are provided to the agent during the simulation.
The operational conditions of the wind turbine are summarized in Table 5. The simulation has been run until the wake has completely developed, reaching the outlet of the domain. In this case, we simulated 30 revolutions of the rotor, around $200\,\mathrm{s}$.
Inflow wind speed [$\mathrm{m/s}$] $11.4$
Rotor speed [$\mathrm{rpm}$] $9.6$
@@
Table 5: Wind turbine DTU 10MW operational conditions.
@@
The solution of the proposed problem is unsteady and a dynamic p-adaptation is performed each time the rotor has rotated $1/16$ of revolution. A snapshot of the average polynomial order and the error estimation, once the wake is fully developed, are represented in Figures 15 and 16. The p-adaptation algorithm has effectively increased the polynomial order to the maximum in the wake and at the BC on the floor. The polynomial order in the rotor plane has been manually limited to a maximum order $p=3$ to fix the modeling errors associated with the actuator line methodology. Also, the far field has been adapted with the minimum order $p=1$, which means that the change in the non-dimensional variables of interest ($\rho u$, $\rho v$ and $\rho w$) is below the defined tolerance (see Section 3.5 for details). Furthermore, downwind turbulent mixing enlarges the wake size, and the high-order region is increased consequently, as shown in Figures 15(a), 15(b), 15(c) and 15(d).
@@
The figures also show the error estimation. The highest error values, around $10^{-2}$, are located near the immersed boundaries and near the floor (no slip boundary condition). Both results were expected, as the mesh was coarsened on purpose to test the effectiveness of the p-adaptation algorithm and the error estimation strategy. Inside the immersed boundaries (tower and nacelle), there is only one h-element of the computational mesh for each height, so even if the polynomial order is high, the error remains large. The interesting point is that the error estimation algorithm has been able to identify the two major sources of error, without having any prior knowledge regarding this specific problem and without training a specific agent for this purpose.
@@ 5 Conclusions
The approach is based on the value iteration algorithm and allows offline training. Therefore, the actual simulations do not see any significant computational overhead. We showcase the methodology for a variety of flows ranging from laminar 2D to turbulent 3D, including circular cylinders, the Taylor Green Vortex, and a 10MW wind turbine. The method offers a robust, reproducible and general approach, applicable to complex three-dimensional problems without needing to retrain. In addition, we include an error estimation, derived from the RL framework, that offers the possibility to quantify local discretization errors.
In general, the findings open up new avenues for further research and potential applications in the solution of complex partial differential equations using RL-based h/p-mesh adaptation.
@@ Acknowledgments
EF and GR acknowledge the funding received by the Grant DeepCFD (Project No. PID2022-137899OB-I00) funded by MCIN/ AEI/10.13039/501100011033 and by ERDF A way of making Europe. OM and EF would like to thank the support of Agencia Estatal de Investigación (for the grant ”Europa Excelencia 2022” Proyecto EUR2022-134041/AEI/10.13039/501100011033) y del Mecanismo de Recuperación y Resiliencia de la Unión Europea. This research has received funding from the European Union (ERC, Off-coustics, project number 101086075). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. Finally, all authors gratefully acknowledge the Universidad Politécnica de Madrid (www.upm.es) for providing computing resources on Magerit Supercomputer and the computer resources at MareNostrum and the technical support port provided by Barcelona Supercomputing Center (RES-IM-2024-1-0003).
@@ Appendix A Compressible Navier-Stokes solver
where $\boldsymbol{q}$ is the state vector of large-scale resolved conservative variables $\boldsymbol{q}=[\rho,\rho v_{1},\rho v_{2},\rho v_{3},\rho e]^{T}$, $\svec{\mathbf{F}}_{e}$ are the inviscid, or Euler fluxes,
$\svec{\mathbf{F}}_{e}=\left[\begin{array}[]{ccc}\rho v_{1}&\rho v_{2}&\rho u_{3}\\ \rho v_{1}^{2}+p&\rho v_{1}v_{2}&\rho v_{1}v_{3}\\ \rho v_{1}v_{2}&\rho v_{2}^{2}+p&\rho v_{2}v_{3}\\ \rho v_{1}v_{3}&\rho v_{2}v_{3}&\rho v_{3}^{2}+p\\ \rho v_{1}H&\rho v_{2}H&\rho v_{3}H\end{array}\right],$ (24)
where $\rho$, $e$, $H=E+p/\rho$, and $p$ are the large-scale density, total energy, total enthalpy and pressure, respectively, and $\vec{v}=[v_{1},v_{2},v_{3}]^{T}$ is the large-scale resolved velocity components. Additionally, $\svec{\mathbf{F}}_{v,turb}$ defines the viscous and turbulent fluxes,

### v2

@@ 4.2 Comparison of p-adaptation approaches and accuracy in lift and drag for a cylinder at Reynolds 100
Once we have tested that the performance of the p-adaptation strategy is appropriate and that we can trust, in general, the error estimation, we increase the complexity of the problem to validate different aspects of the RL p-adaptation approach applied to the Navier-Stokes equations. In particular, to improve the quality of the designed test cases and quantitatively measure the performance of the p-adaptation strategy, we compare our novel approach with a p-adaptation sensor based on the original sensor developed by Persson and Peraire [19]. This sensor compares the solution inside each element using a different number of high-order modes, providing a spectral decay indicator that can be used as an estimation of the smoothness of the solution. Although it was first developed for shock-capturing purposes, based on this sensor some p-adaptation approaches have been successfully used in DG solvers [17]. However, the available strategies based on this sensor are very limited, as they have only been used to increase the polynomial order (not to decrease it), and they cannot handle anisotropic adaptation. We have made slight improvements in the sensor-based approaches found in the literature by incorporating these two additional functionalities. In this way, we ensure a fair comparison between our novel method and the state-of-the-art sensor.
The problem to be solved is the laminar, but unsteady, flow at $Re=100$ around a cylinder at $M=0.15$. The mesh used for this problem is shown in Figure 10. This coarse mesh has only 1282 elements on purpose, again to highlight the ability of the RL agent to increase the polynomial where required. However, the elements are slightly smaller downstream to capture the vortex shedding.
@@
Figure 10: Representation of the mesh used for the simulation of the flow around a cylinder at $Re=100$ and $M=0.15$.
For this problem, a dynamic p-adaptation is required, as the solution is unsteady. A snapshot of the average polynomial order in each element using both approaches (RL and Persson and Peraire-based sensor), once the flow field is stabilized, is represented in Figure 11. At first glance, both approaches are similar, as they increase the polynomial near the cylinder and in the wake, where the vortex shedding is present. However, the adaptation performed by our RL agent is more efficient, as it increases the polynomial order only where required, while the Persson and Peraire-based approach increases the polynomial order in some regions upstream and also downstream far from the wake, where the flow field is almost steady.
It is very important to mention that, on the one hand, the Persson and Peraire-based approach relies on a threshold to decide when to increase and decrease the polynomial order, and this threshold is problem-dependent. It took several attempts of trial and error, using different values of the aforementioned threshold, to achieve the polynomial adaptation shown in Figure 11(b). On the other hand, the parameters of our RL agent are decided beforehand, before the training, and their dependency with each specific problem is small. Therefore, our agent does not require problem-specific tuning, which is an important advantage of the proposed methodology, when compared to other state-of-the-art p-adaptation strategies.
For completeness, a snapshot of the velocity field is represented in Figure 12. The vortex shedding is well captured, as expected. Finally, we report the accuracy of the p-adaptation approach in Table 3. A comparison between the average drag, the maximum lift and the Strouhal number is considered. The results obtained with the proposed methodology are very close to those of other sources from the scientific literature.
@@
Figure 11: Comparison of p-adaptation between a) the RL agent strategy and b) the improved Persson and Peraire sensor for a simulation of the flow around a cylinder at $Re=100$ and $M=0.15$. A snapshot of the average polynomial order in each element of the mesh is illustrated.
@@
Figure 12: Snapshot of the velocity field for the simulation of the flow around a cylinder at $Re=100$ and $M=0.15$ using the RL p-adaptation agent. Both, a) the horizontal velocity and b) the vertical velocity are shown.
@@ 4.3 Laminar - turbulent transition for the Taylor Green Vortex problem at Reynolds 1600
The flow field is continuously evolving and it has been widely reported that under-resolution without special numerical treatment creates poor results [57, 58, 59, 60]; therefore, a dynamic p-adaptation is an excellent option to accurately solve the problem while maintaining the cost at minimum. A p-adaptation is performed at every $0.05$ non-dimensional time units. The non-dimensional velocity magnitude and the polynomial adaptation for different timestamps are shown in Figure 13. Each snapshot shows how the p-adaptation algorithm is able to capture velocity fluctuations by increasing the polynomial order locally. As time progresses, the flow transitions to a turbulent state, and hence the average polynomial order in the whole domain increases. This phenomenon is highlighted in Figure 14(a), where it is clear how the polynomial order is progressively increased, almost reaching the maximum polynomial order $p_{\max}=6$ after 20 non-dimensional time units (fully turbulent regime).
Finally, in Figure 14(b) we show the evolution of the kinetic energy rate over time for different solutions: two uniform polynomial orders $p=3$ and $p=6$, the RL p-adaptation and a DNS solution from [61]. On the one hand, the coarse $p=3$ solution presents large deviations from the reference DNS solution at the transitional times $t/t_{c}\in[8,14]$. On the other hand, both the p-adapted and the uniform $p=6$ solution are very close to the reference DNS simulation, with both matching almost perfectly. Despite these similar results between the $p=6$ and the p-adapted solutions, the simulation with dynamic p-adaptation was much faster, as described in Table 4. The novel p-adaptation strategy is shown to be accurate in the presence of turbulence while significantly reducing the computational cost by 60%.
@@ 4.4 Simulation of the wind turbine DTU 10MW
Finally, once the reinforcement learning agent has been tested in a wide variety of scenarios and problems, we apply this methodology to solve the flow field around a wind turbine. The objective behind this problem is to demonstrate that the proposed p-adaptation approach can be used in complex scenarios without any previous knowledge regarding the simulation; that is, it is able to adapt to different situations even though it has not been trained for any specific case.
We focus on the 10 MW DTU wind turbine, whose blades are modeled using an actuator line [62, 63, 64], while the nacelle and pylon have been simulated using an immersed boundary approach [65, 49]. Furthermore, a no-slip boundary condition has been set at the bottom of the domain, while free-slip boundary conditions have been applied at the top and at the side faces of the domain. Each of these elements adds an additional level of complexity to the p-adaptation algorithm:
@@
    Behind the rotor, in the wake of the wind turbine, the flow is fully turbulent, and high polynomials are required. An accurate computation of this region is essential to obtain accurate predictions of the power generated by the wind turbine.
@@
    An LES model is used to simulate under-resolved scales. In this case, we use the Vreman model [66], detailed in A.
The aforementioned elements must be taken into account to obtain an accurate solution. In our case, we use a coarse and almost uniform Cartesian h-mesh with $294492$ hexahedral elements, and the RL agent will select the appropriate polynomial order inside each element. We limit the maximum polynomial order to $p=4$ in the domain and to $p=3$ in the rotor plane, to fix modeling errors associated with the actuator line methodology. Again, it is exactly the same RL agent that was used in all the previous cases and it has no knowledge in relation to this specific problem. Despite this, we expect to obtain an appropriate polynomial adaptation based on the values at the Gauss nodes (the state) that are provided to the agent during the simulation.
The operating conditions of the wind turbine are summarized in Table 5. The simulation has been run until the wake has completely developed, reaching the outlet of the domain. In this case, we simulated 30 revolutions of the rotor, around $225\,\mathrm{s}$.
Inflow wind speed [$\mathrm{m/s}$] $10.0$
Rotor speed [$\mathrm{rpm}$] $8.06$
@@
Table 5: Wind turbine DTU 10MW operating conditions.
The solution is unsteady and turbulent, and a dynamic p-adaptation is performed each time the rotor rotates $1/16\mathrm{th}$ of revolution. First, we validate the results with published data and explore the proposed error estimation. To do so, an additional simulation of the proposed configuration and operational conditions has been carried out, only this time allowing a maximum polynomial order of $p=6$ in the entire domain and setting a constant polynomial order of $p=5$ in the rotor zone. Figure 15 shows a comparison between the axial and tangential force distributions along the blade span, given by both simulations and the blade-resolved simulation presented in [67], which corresponds to the same set of operational conditions. In this case, the increase in the maximum allowed polynomial order results only in a slight improvement of the solution with respect to the blade-resolved simulation, which overall shows good agreement with the exception of the axial components near the tip of the blade. These small differences are expected in actuator line simulations, as the source term used to represent rotating blades is calculated using 2D airfoil data (including a tip correction), but the flow has additional 3D effects at the root and tip of the blades [68, 64]. Figure 16(a) compares the axial velocity of the wake averaged in time (over three rotations), at the hub height and a distance of 2.3D downstream of the rotor plane, where D is the diameter of the rotor, between the numerical simulations and the experimental results presented in [69], which were obtained in a wind tunnel for a model of scale $1/75\mathrm{th}$ of the DTU 10MW wind turbine. Simulation results provide a reasonable representation of the characteristic double Gaussian distribution found in the experimental setup, whereas again the increase in the maximum polynomial order does not translate into a significant improvement in results, illustrating the correctness of the initial simulation. The absolute difference between the time-averaged wake velocities of the simulations with maximum polynomial orders of $p=4$ and $p=6$ is presented in Figure 16(b), as a measure of the error of the first simulation setup. This distribution is then used to assess the performance of the proposed RL-based error estimator (maximum error) computed for the simulation in which $p=4$ is used as the maximum polynomial order, which provides a good prediction of the shape, trend, and magnitude of the error, even capturing the drop at approximately $y/D=0.3$ and the peak at $y/D=0.6$. This estimation provides valuable insight into the error, as it can be computed inexpensively for any simulation without prior knowledge of the solution.
Figure 15: Comparison of (a) axial and (b) tangential force distributions along rotor span between blade-resolved RANS simulation in [67] (blue) and HORSES3D LES + actuator lines with polynomial orders of $p=3$ and $p=5$ in the rotor zone (orange and green, respectively).
Figure 16: Axial velocity of the wake averaged in time (over three rotations), at the hub height and a distance of 2.3D downstream of the rotor plane, where D is the diameter of the rotor. Comparison between (a) experimental results presented in [69] (blue) and HORSES3D LES + actuator line simulation with maximum polynomial orders of $p=4$ and $p=6$ (orange and green, respectively) and (b) absolute error in time-averaged axial velocity of the wake for a max. $p=4$ with respect to a max. $p=6$ simulation (blue) and RL-based error estimation (maximum error) for a max. $p=4$ simulation (orange).
@@
For completeness, a snapshot of the average polynomial order and the error estimation, once the wake has fully developed, are represented in Figures 17 and 18. The p-adaptation algorithm has effectively increased the polynomial order to the maximum in the wake and at the BC on the floor. Also, the agent has selected the minimum polynomial order $p=1$ for the far field region, which means that the change in the non-dimensional variables of interest ($\rho u$, $\rho v$ and $\rho w$) is below the defined tolerance (see Section 3.5 for details). Furthermore, downwind turbulent mixing enlarges the wake size, and consequently the high-order region is increased, as shown in Figures 17(a), 17(b), 17(c) and 17(d).
@@
Finally, these figures also show the error estimation for the non-dimensional axial velocity at different locations. The highest error values, around $10^{-2}$, are located near the immersed boundaries and near the floor (no slip boundary condition). Both results were expected, as the mesh was coarsened on purpose to test the effectiveness of the p-adaptation algorithm and the error estimation. Inside the immersed boundaries (tower and nacelle), there is only one h-element of the computational mesh for each height, so even if the polynomial order is large, the error remains large. The error estimation algorithm is able to identify the two major sources of error, without having any prior knowledge regarding this specific problem and without training a specific agent for this purpose. Furthermore, the computational cost of the estimator is negligible, making it a good choice to capture the trend of spatial errors in numerical simulations.
@@ 5 Conclusions
The approach is based on the value iteration algorithm and allows offline training, which is decoupled from the CFD solver, leading to a flexible and efficient approach. Therefore, the actual simulations do not show any significant computational overhead. We showcase the methodology for a variety of flows ranging from laminar 2D to turbulent 3D, including circular cylinders, the Taylor Green Vortex, and a 10MW wind turbine. The method offers a robust, reproducible, and general approach, applicable to complex three-dimensional problems without needing to retrain. In addition, we include an inexpensive error estimation, derived from the RL framework, that offers the possibility to quantify local discretization errors with a negligible cost in numerical simulations.
These findings open new avenues for further research and potential applications in the solution of complex partial differential equations using RL-based h/p-mesh adaptation.
@@ Acknowledgments
This research has received funding from the European Union (ERC, Off-coustics, project number 101086075). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. EF and GR acknowledge the funding received by the Grant DeepCFD (Project No. PID2022-137899OB-I00) funded by MICIU/AEI/10.13039/501100011033 and by ERDF, EU. OM and EF would like to thank the support of Agencia Estatal de Investigación (for the grant ”Europa Excelencia 2022” Proyecto EUR2022-134041/AEI/10.13039/501100011033) y del Mecanismo de Recuperación y Resiliencia de la Unión Europea. DH and EF acknowledge the funding received by the Comunidad de Madrid according to Orden 5067/2023, of December 27th, issued by the Consejero de Educación, Ciencia y Universidades, which announces grants for the hiring of predoctoral research personnel in training for the year 2023. Finally, all authors gratefully acknowledge the Universidad Politécnica de Madrid (www.upm.es) for providing computing resources on Magerit Supercomputer and the computer resources at MareNostrum and the technical support provided by Barcelona Supercomputing Center (RES-IM-2024-1-0003).
@@ Appendix A Compressible Navier-Stokes solver
where $\boldsymbol{q}$ is the state vector of large-scale resolved conservative variables $\boldsymbol{q}=[\rho,\rho u,\rho v,\rho w,\rho e]^{T}$, $\svec{\mathbf{F}}_{e}$ are the inviscid, or Euler fluxes,
$\svec{\mathbf{F}}_{e}=\left[\begin{array}[]{ccc}\rho u&\rho v&\rho w\\ \rho u^{2}+p&\rho uv&\rho uw\\ \rho uv&\rho v^{2}+p&\rho vw\\ \rho uw&\rho vw&\rho w^{2}+p\\ \rho uH&\rho vH&\rho wH\end{array}\right],$ (24)
where $\rho$, $e$, $H=E+p/\rho$, and $p$ are the large-scale density, total energy, total enthalpy and pressure, respectively, and $\vec{v}=[u,v,w]^{T}$ is the large-scale resolved velocity components. Additionally, $\svec{\mathbf{F}}_{v,turb}$ defines the viscous and turbulent fluxes,